OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (anonymous):

is that cos4x or cos^4(x) and is it cos2x or cos^2(x) @JoshKPatterson

OpenStudy (anonymous):

cos 4x + cos 2x. No exponents on them.

OpenStudy (anonymous):

@genius12

OpenStudy (mertsj):

\[\cos 4x+\cos 2x=\cos (2x+2x)+\cos 2x=\cos 2x \cos 2x-\sin 2x \sin 2x+\cos 2x\]

OpenStudy (mertsj):

\[(2\cos ^2-1)(1-2\sin ^2)-2\sin x \cos x2\sin x \cos x+2\cos ^2x-1\]

OpenStudy (mertsj):

\[2\cos ^2x-4\sin^2x \cos ^2x-1+2\sin ^2x-4\sin ^2x \cos ^2x+2\cos ^2x-1\]

OpenStudy (mertsj):

\[4\cos ^2x-8\sin ^2x \cos ^2x+2\sin ^2x-2\]